MEDGAR EVERS COLLEGE he City University of New York Basic Physics PHY 114 - W-01 Department of Physical and Computer Sciences Medgar Evers College Assignment 2 Name: Write or type work and solutions on a separate paper. If written, write legibly. 1. Find the magnitude and direction (polar coordinates) of the difference vector 20 m due east and B 10 m due north. 2. Find the vector components (rectangular coordinates) of 2A +v3B where 3. Use the graphical tail-to-head technique to find the vector sums of the vectors B A where A A 4m/s2 at 60° west of south and B -3m/s at 60 south of west. A a velocity of 20 m/s at 30° north of east B a velocity of 50 m/s at 45 north of east 4. Extra Credit: A hiker begins a trip by first walking 25.0 km southeast from her car. She stops and sets up her tent for the night. On the second day, she walks 40.0 km in a direction 60.0 north of east, at which point she discovers a forest rangers tower. If the hiker wishes to return to her car from the rangers tower along a single straight line what are the components of the vector representing this hike and it cardinal direction?
Answer #1

1.

assuming east-west direction along x-axis and north-south direction along y-axis

A = 20 i

B = 10 j

difference is given as

B - A = 10 j - 20 i

B - A = - 20 i + 10 j

magnitude using Pythagorean theorem is given as

|B - A| = sqrt((- 20)2 + (10)2) = 22.4 m

direction : 180 - tan-1(10/20) = 153.4 deg counterclockwise from positive x-axis